Calculate the income elasticities for the following demand functions.


a) x(p,m) = m/2p*



b) x(p,m) = 10/p*


1
Expert's answer
2021-11-24T14:22:31-0500

Demand is given by "x =\frac{ m}{(2p)}"

Price elasticity is given by

"Ep=\frac{ \frac{dx}{x}}{ \frac{dp}{p}} =\frac{ dx}{dp} \times \frac{p}{x}"

"\frac{dx}{dp} = \frac{-m}{2} \times \frac{ 1}{p\\^2}"

Plugging in,

"Ep=\frac{ -m}{2} \times \frac{ 1}{p^2} \times \frac{2p^2}{m} = -1"

Income elasticity is given by

"Em =\frac{\frac{dx}{x}}{ \frac{dm}{m}} = \frac{dx}{dm} \times \frac{m}{x}"

"\frac{dx}{dm} =\frac{ 1}{2p}"

Plugging in,

"Em= \frac{1}{(2p)} \times 2p = 1"